OpenStudy (anonymous):

Solve 2x + 5y = –13 3x – 4y = –8

OpenStudy (anonymous):

solve for both x first then y and then plug them back into the problem for the final answer

OpenStudy (anonymous):

x=-4 y=-1

OpenStudy (anonymous):

yea now plug them in to the original problem and then get ur answer

OpenStudy (anonymous):

yeah right bodacious

OpenStudy (anonymous):

that is how i would doo it

OpenStudy (anonymous):

here is the solving your problem as attached

OpenStudy (anonymous):

oh never mind ur doing something totally different sorry

OpenStudy (anonymous):

2x + 5y = –13 2x+5y=-13 Since 2x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2x from both sides. 5y=-2x-13 Divide each term in the equation by 5. 5y 5 =-2x 5 -13 5 Simplify the left-hand side of the equation by canceling the common terms. y=-2x 5 -13 5 Combine the numerators of all expressions that have common denominators. y= -2x-13 5

OpenStudy (anonymous):

3x-4y=-8 Since -4y does not contain the variable to solve for, move it to the right-hand side of the equation by adding 4y to both sides. 3x=4y-8 Divide each term in the equation by 3. 3x 3 = 4y 3 -8 3 Simplify the left-hand side of the equation by canceling the common terms. x= 4y 3 -8 3 Simplify the right-hand side of the equation by simplifying each term. x= 4(y-2) 3